Find the volume of the region that lies above the cone z= sqrt(x^2+y^2) and within the unit sphere x^2+y^2+z^2= 1 using spherical coordinates
1
Expert's answer
2020-11-12T18:52:15-0500

The region of integration is a cone in spherical coordinates where radius changes from 0 to 1,

azimuthal angle "\phi" changes from 0 to "2\pi" and polar angle "\theta" changes from 0 to"\frac{\pi}{4}" .


"V=\iiint_{V}{dV}=\int_{0}^{2\pi}\int_{0}^{\frac{\pi}{4}}\int_{0}^{1}r^2\sin{\theta} dr d\theta d\phi="


"=\int_{0}^{1}r^2dr \int_{0}^{\frac{\pi}{4}} \sin\theta d\theta \int_{0}^{2\pi}d\phi" =


"=\frac{r^3}{3}|_{0}^{1}* (-\cos{\theta})|_{0}^{\frac{\pi}{4}}*\phi|_{0}^{2\pi}="


"=\frac{1}{3}(-\frac{\sqrt{2}}{2}-(-1))*2\pi="


"=\frac{\pi}{3}(2-\sqrt{2})"